OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (4, -6) m = 3/5

OpenStudy (anonymous):

SO many replys :P

OpenStudy (anonymous):

y = 3/5x + m -6 = 3/5 (4) +m Solve for m, plug back into first equation.

OpenStudy (anonymous):

point slope form is like: y-y1 = m(x-x1) y1 = -6, x1 = 4, m = 3/5

OpenStudy (anonymous):

so you get: y+6=(3/5)(x-4)